The square of an integer plus two times the integer is 35. Find all possible solutions.
Answer by mananth(16949)

let integer be x
x^2 +2x=35
x^2+2x-35=0
x^2+7x-5x-35=0
x(x+7)-5(x+7)=0
(x+7)(x-5)=0
x=-7 or 5